# On the identifiability of causal graphs with multiple environments

### v1

@@ 1 Introduction
Causal discovery seeks to recover cause–effect structure from data, which allows counterfactual reasoning and prediction under interventions (Pearl, 2009; Peters et al., 2017; Spirtes, 2010; Spirtes et al., 2000). However, learning causal structure from purely observational i.i.d. data is, in general, ill-posed: multiple directed acyclic graphs (DAGs) are distributionally equivalent, i.e., indistinguishable from the data distribution.
In the interventional causal discovery literature, hard interventions—directly modifying the causal structure—are known to unlock identifiability of the underlying graph. Classic results from Eberhardt et al. (2005) show that the number of sufficient hard interventions to identify the causal order scales logarithmically with the number of nodes. A large corpus of intervention-based causal discovery research has largely built on these findings (Eberhardt, 2008; He & Geng, 2008; Hauser & Bühlmann, 2012; Shanmugam et al., 2015; Kocaoglu et al., 2017; Wang et al., 2017; Lindgren et al., 2018; Eaton & Murphy, 2007; Triantafillou & Tsamardinos, 2015; Lorch et al., 2022; Ke et al., 2023b).
Recent work has explored the problem of causal graph identifiability from multiple environments and soft interventions (i.e., in the setting where non i.i.d. data might naturally occur and does not stem from changes in the causal structure) (Perry et al., 2022; Huang et al., 2020; Heinze-Deml et al., 2018; Peters et al., 2015; Ghassami et al., 2017, 2018; Jaber et al., 2020; Jalaldoust et al., 2025; Brouillard et al., 2020; Heurtebise et al., 2025): however, from an identifiability perspective, these results do not provide guarantees of recovery of the unique causal graph with a limited number of environments under generic assumptions.
@@
Our work is also of independent methodological interest. In particular, key to our theory is the duality between causal discovery and independent component analysis (ICA). Reizinger et al. (2023) recently formalized that nonlinear ICA identifiability results naturally extend to structure learning (well known in the linear case since Shimizu et al. (2006)). This is of great relevance in light of the late advancements in multi-environment ICA identifiability pioneered by Hyvärinen & Morioka (2016); however, directly bootstrapping these findings to causal discovery doesn’t carry great promise, being ICA the harder problem of the two: we show that where ICA identifiability requires a number of environments that scales linearly with the number of variables, causal graph identifiability can be achieved with data from just two extra domains. This calls for causality-only identifiability results in the multi-environment setting, as developed in our work. Inspired by the recent success of ICA with multiple environments, we are hopeful that our approach paves the way to novel causality theory that weakens the requirements in terms of heterogeneity of the data and parametric assumptions.
@@ 2 Related works
Several works in the literature have addressed causal discovery identifiability and estimation via non i.i.d. data (stemming from soft interventions and multiple environments). Peters et al. (2015) and Heinze-Deml et al. (2018) identify the parents of a designated target node via invariance across environments, yielding partial identifiability of causal directions. Huang et al. (2020) use nonstationarity to recover the skeleton and orient some edges. Perry et al. (2022) leverage sparse mechanism shifts, proving high-probability graph recovery with bounds that improve as the number of environments grows. Ghassami et al. (2017, 2018) and Heurtebise et al. (2025), similarly to our work, study identifiability of structural causal models from multiple environments, but their identifiability results are specialized to the linear case. Recently, Jalaldoust et al. (2025) formulated a statistical test that can find a superset of the parents of a target node. Yang et al. (2018); Brouillard et al. (2020); Jaber et al. (2020) characterize equivalence classes identifiability from interventions. From a methodological perspective, Brouillard et al. (2020); Ke et al. (2023a) introduce differentiable approaches to causal discovery with interventions; Mooij et al. (2020) propose a unifying framework for causal discovery from observational and multi-environment data. All of these results are complementary to our work, which is, to the best of our knowledge, the first to provide guarantees of identifiability of the causal graph from a finite number of auxiliary additional environments, potentially only two.
@@
The seminal work of Shimizu et al. (2006) shows that if an SCM can be expressed as a linear non-Gaussian ICA model, the underlying causal graph is identifiable. Reizinger et al. (2023) generalize this to the nonlinear case. Monti et al. (2020) show that time contrastive ICA (Hyvärinen & Morioka, 2016) can identify bivariate causal graphs with arbitrary nonlinear mechanism. The common ground of these findings is that they adapt the existing ICA identifiability theory to the problem of causal discovery. This approach is clearly important, especially in the light of the recent advancement in multi-environment ICA identifiability (Hyvärinen et al., 2019; Khemakhem et al., 2020a, b; Gresele et al., 2019; Hälvä & Hyvärinen, 2020; Hyvärinen & Morioka, 2017; Hälvä et al., 2021); however, in the nonlinear setting, it fails to capture the gap between the two problems: while ICA attempts to recover the mixing function and the independent sources at each point, causal discovery concerns the much simpler problem of structure identifiability. Our work shows that this difference is key to demonstrating causal discovery identifiability from a constant number of sufficiently different environments, where ICA requires at least as many as the number of sources (see e.g. Theorem 1 in Hyvärinen & Morioka (2016)).
@@ 3.1 Structural causal models and ICA
It is known (Reizinger et al., 2023) that, under some faithfulness assumption, the support of the Jacobian of the mixing function completely identifies the causal structure.
@@
Proposition 1 (Proposition 1 in Reizinger et al. (2023)).
@@
This formulation of faithfulness is well known and at the core of the LiNGAM algorithm for linear SCMs (Shimizu et al., 2006), and is satisfied almost everywhere under some regularity conditions on $\mathbf{f}$. When this is the case, the above proposition means that for causal discovery we are interested in the support of the inverse Jacobian, and, by Equation˜3, this can be recovered by having access to the support at a single point where faithfulness is satisfied.
@@ 4 Theory
˜1 is standard when proving identifiability: the results in Hoyer et al. (2008); Zhang & Hyvärinen (2009); Immer et al. (2022) are based on higher-order derivatives, and have strong requirements that guarantee invertibility. Also ˜2 is mild and somewhat necessary: it simply asks that the interventions are meaningful, i.e. that they affect the variance; interventions on the mean, intuitively, are not informative as they shift the density graph by a constant, without affecting its shape (the gradient and the Hessian of the density, where information about the causal graph lies). ˜3 requires that the Jacobian of the inverse of the mixing function is informative about the causal structure at the mean of $\mathbf{S}$ (and it’s almost surely verified over $\mathbf{X}$ samples, under some generic regularity conditions on $\mathbf{f}$). The reason behind it is that we probe the identifiability of the Jacobian’s support at the mean. The only real simplifying constraint is ˜4 of the Gaussianity of the sources, which is, however, not new in the literature (see e.g.Rolland et al. (2022)). Later, we discuss why this assumption is needed in the paper and potential ways to relax it.
@@ 4.1 Identifiability from second order derivatives of the log-likelihood
The assumption basically requires that there is sufficient variability between the different environments. Similar requirements of sufficient variability are ubiquitous in the nonlinear ICA literature (e.g. Hyvärinen & Morioka (2016); Khemakhem et al. (2020b); Lachapelle et al. (2022)). Intuitively speaking, ˜5 is satisfied when, for each of the two groups of environments ($1,...,e_{1}$ and $e_{1}+1,...,e_{1}+e_{2}$), each source $S_{j}$ is subject to rescaling. To see that, consider the LHS of the first equation: $\lambda_{j}^{i}=1$ for each $i=1,...,e_{1}$ corresponds to the case when the variable $S_{j}$ is never subject to rescaling in any of the environments, and indeed yields a violation of the assumption. Note that even if $S_{j}$ is subject to rescaling for some index $i$, the values of $(\lambda_{j}^{i})_{i\in[e_{1}]}$ can always be tuned such that the assumption is violated; however, this corresponds to pathological choices of the rescaling coefficients, which never occur in general (shown in Proposition˜3 in the appendix).
@@ 5 Empirical results
In this section, we report and analyse empirical results that validate our theory. Our experiments on synthetic data show that if the assumptions of Theorem˜1 hold, the causal direction can be recovered from the data. We focus on bivariate graphs, commonly adopted as the easiest yet non-trivial setting for testing identifiability (e.g., Hoyer et al. (2008); Zhang & Hyvärinen (2009); Immer et al. (2022)).
@@ 5.1 Synthetic data generation
We generate synthetic data from bivariate causal models with independent noise terms, sampled from a normal distribution with unit mean and covariance entries uniformly drawn between $[1,1.5]$. Given the variables $x_{1},x_{2}$ and the graph $x_{1}\to x_{2}$ we consider the following causal mechanisms that comply with the assumptions of Theorem˜1: (i) $x_{2}:=s_{1}^{2}\operatorname{arctan}(s_{2})+s_{2}^{3}$ (ii) $x_{2}:=s_{1}^{2}s_{2}+\operatorname{arctan}(s_{2})$ (iii) $x_{2}:=s_{1}^{2}+\operatorname{arctan}(s_{1})s_{2}+s_{1}s_{2}^{3}$. Note that any of these models can not be reparametrized to a post nonlinear or location scale noise model, which are the most general SCMs identifiable from pure observations (Zhang & Hyvärinen, 2009; Immer et al., 2022). Additionally, we consider data from a linear Gaussian model, notably non-identifiable. We run experiments on datasets with $\{3,6,9\}$ environments. For each environment, we generate $2000$ observations. In Section˜C.4, we discuss experiments with non-Gaussian independent sources. Interestingly, these additional results seem to support our hypothesis that Theorem˜1 could be extended to other source distributions.
@@ 5.2 Analysis of the experimental results
The simplified pseudocode is found in Algorithm˜1 (a detailed version is presented in Section˜C.3). The steps in our procedure closely follow the proof of Theorem˜1: this approach to algorithmic design is not necessarily the best, which is why we highlight that our method is not within our main contributions. For a single inference, the input is the data tensor $\widehat{X}\in\mathbb{R}^{k\times n\times d}$: for each environment from $1$ to $k$ it consists of a dataset with $n$ observations of $d$ causal variables. Additionally, we are given the sets $I_{1},I_{2}\subset[k]$ of indices that split the auxiliary environments into two groups, as required by our theory. The first environment is taken as the base one. We have two steps where statistical estimation is involved: (i) For each environment, the gradient and the Hessian of the log-likelihood are approximated via the Stein gradient estimator, introduced in Li & Turner (2018) and popularized in causal discovery by Rolland et al. (2022); Montagna et al. (2023b); (ii) For each environment $i\in[k]$, we need to find the observation $j\in[n]$ such that $\mathbf{f}^{-1}(\widehat{X}[i,j])\approx\mu_{\mathbf{S}}$, that is, the data point generated mixing the source vector at the mean. Fortunately, this can be consistently estimated from the score $\nabla\log p_{\mathbf{x}}$, as we demonstrate in Proposition˜2 in the appendix. These two steps are achieved by Algorithm˜1 at the end of the first for loop. At this stage, all statistical quantities have been estimated: we note that, being the Stein estimator consistent, the algorithm is correct in the infinite sample limit. In the second for loop, we take the points at the estimated mean that we previously found, and compute the difference of the Hessians between the base and auxiliary environments: this exactly mirrors the first equality in Equations˜9 and 10 of Lemma˜1. Next, in the algorithm’s notation, we compute
@@
Then, we solve the linear system $\widehat{H}_{\textnormal{diffs}}[1]M=\widehat{H}_{\textnormal{diffs}}[2]$ to find $M$. In the infinite samples limit Equation˜13 is a precise equality, such that $M$ and $\Omega_{1}^{-1}\Omega_{2}$ are similar: diagonalizing $M$ we find $J_{\mathbf{f}^{-1}}$ up to a scaled permutation. The permutation indeterminacy is removed leveraging the assumption that the causal graph is acyclic via standard arguments (see Shimizu et al. (2006) and Reizinger et al. (2023)). Finally, the algorithm returns the estimated support of the inverse Jacobian.
@@ A.4 Proof of Theorem 1
By Lemma 1 in Reizinger et al. (2023), the permutation indeterminacy can be uniquely determined and thus removed. Given that by ˜3 the Jacobian of $J_{\mathbf{f}^{-1}}(\mathbf{x})$ is faithful to the causal graph, the claim is verified.
@@ Appendix B Independent component analysis
In this section, we present a primer on the problem of Independent Component Analysis (ICA), based on the content of Section 2 in Buchholz et al. (2022). ICA seeks to recover latent sources from their observed mixtures. We assume a hidden random vector $\mathbf{S}\in\mathbb{R}^{d}$ with independent coordinates and observations generated by
@@
In general (i.e., for $(\mathcal{F},\mathcal{P})$ arbitrarily large), the ICA problem is not identifiable for reasonable $\mathcal{S}$. Notable example comes from the Darmois construction or constructions based on measure-preserving transformations. Several results in the literature have studied which conditions on $(\mathcal{F},\mathcal{P})$ can help identifiability. Most notably, Buchholz et al. (2022) shows that when $\mathcal{F}$ represents the class of conformal maps, identifiability is guaranteed up to trivial indeterminacies. If heterogeneous data are considered (e.g., in the multi-environment setting of this paper), identifiability was shown in the general case (Hyvärinen & Morioka, 2016).

### v2

@@ 1 Introduction
Causal discovery seeks to recover cause–effect structure from data, which allows counterfactual reasoning and prediction under interventions (pearl2009causality; peters2017elements; spirtes2010introduction; Spirtes2000). However, learning causal structure from purely observational i.i.d. data is, in general, ill-posed: multiple directed acyclic graphs (DAGs) are distributionally equivalent, i.e., indistinguishable from the data distribution.
In the interventional causal discovery literature, hard interventions—directly modifying the causal structure—are known to unlock identifiability of the underlying graph. Classic results from eberhardt2005on show that the number of sufficient hard interventions to identify the causal order scales logarithmically with the number of nodes. A large corpus of intervention-based causal discovery research has largely built on these findings (eberhardt2008almost; he2008active; hauser2012characterization; shanmugam2015learning; kocaoglu2017costoptimal; wang2017permutationbased; lindgren2018experimental; eaton2007exact; triantafillou2015constraintbased; lorch2022amortized; ke2023learning).
Recent work has explored the problem of causal graph identifiability from multiple environments and soft interventions (i.e., in the setting where non i.i.d. data might naturally occur and does not stem from changes in the causal structure) (perry2022causal; huang2020heterogeneous; heinze2018invariant; peters2015invariant; ghassami2017learning; ghassami2018multidomain; jaber2020causal; jalaldoust2025multidomain; brouillard2020differentiable; heurtebise2025identifiablemultiviewcausaldiscovery): however, from an identifiability perspective, these results do not provide guarantees of recovery of the unique causal graph with a limited number of environments under generic assumptions.
@@
Our work is also of independent methodological interest. In particular, key to our theory is the duality between causal discovery and independent component analysis (ICA). reizinger2023jacobianbased recently formalized that nonlinear ICA identifiability results naturally extend to structure learning (well known in the linear case since shimizu2006alinear). This is of great relevance in light of the late advancements in multi-environment ICA identifiability pioneered by hyvarinen2016time; however, directly bootstrapping these findings to causal discovery doesn’t carry great promise, being ICA the harder problem of the two: we show that where ICA identifiability requires a number of environments that scales linearly with the number of variables, causal graph identifiability can be achieved with data from just two extra domains. This calls for causality-only identifiability results in the multi-environment setting, as developed in our work. Inspired by the recent success of ICA with multiple environments, we are hopeful that our approach paves the way to novel causality theory that weakens the requirements in terms of heterogeneity of the data and parametric assumptions.
@@ 2 Related works
Several works in the literature have addressed causal discovery identifiability and estimation via non i.i.d. data (stemming from soft interventions and multiple environments). peters2015invariant and heinze2018invariant identify the parents of a designated target node via invariance across environments, yielding partial identifiability of causal directions. huang2020heterogeneous use nonstationarity to recover the skeleton and orient some edges. perry2022causal leverage sparse mechanism shifts, proving high-probability graph recovery with bounds that improve as the number of environments grows. ghassami2017learning; ghassami2018multidomain and heurtebise2025identifiablemultiviewcausaldiscovery, similarly to our work, study identifiability of structural causal models from multiple environments, but their identifiability results are specialized to the linear case. Recently, jalaldoust2025multidomain formulated a statistical test that can find a superset of the parents of a target node. yang2018characterizing; brouillard2020differentiable; jaber2020causal characterize equivalence classes identifiability from interventions. From a methodological perspective, brouillard2020differentiable; ke2023neural introduce differentiable approaches to causal discovery with interventions; mooij2020jci propose a unifying framework for causal discovery from observational and multi-environment data. All of these results are complementary to our work, which is, to the best of our knowledge, the first to provide guarantees of identifiability of the causal graph from a finite number of auxiliary additional environments, potentially only two.
@@
The seminal work of shimizu2006alinear shows that if an SCM can be expressed as a linear non-Gaussian ICA model, the underlying causal graph is identifiable. reizinger2023jacobianbased generalize this to the nonlinear case. monti2019nonsens show that time contrastive ICA (hyvarinen2016time) can identify bivariate causal graphs with arbitrary nonlinear mechanism. The common ground of these findings is that they adapt the existing ICA identifiability theory to the problem of causal discovery. This approach is clearly important, especially in the light of the recent advancement in multi-environment ICA identifiability (hyvarinen2019nonlinear; khemakhem2020variational; khemakhem2020icebeem; gresele2019incomplete; halva2020hidden; hyvarinen2017nonlinear; halva2021snica); however, in the nonlinear setting, it fails to capture the gap between the two problems: while ICA attempts to recover the mixing function and the independent sources at each point, causal discovery concerns the much simpler problem of structure identifiability. Our work shows that this difference is key to demonstrating causal discovery identifiability from a constant number of sufficiently different environments, where ICA requires at least as many as the number of sources (see e.g. Theorem 1 in hyvarinen2016time).
@@ 3.1 Structural causal models and ICA
It is known (reizinger2023jacobianbased) that, under some faithfulness assumption, the support of the Jacobian of the mixing function completely identifies the causal structure.
@@
Proposition 1 (Proposition 1 in reizinger2023jacobianbased).
@@
This formulation of faithfulness is well known and at the core of the LiNGAM algorithm for linear SCMs (shimizu2006alinear), and is satisfied almost everywhere under some regularity conditions on $\mathbf{f}$. When this is the case, the above proposition means that for causal discovery we are interested in the support of the inverse Jacobian, and, by Equation˜3, this can be recovered by having access to the support at a single point where faithfulness is satisfied.
@@ 4 Theory
˜1 is standard when proving identifiability: the results in hoyer2008anm; zhang2009pnl; immer2022on are based on higher-order derivatives, and have strong requirements that guarantee invertibility. Also ˜2 is mild and somewhat necessary: it simply asks that the interventions are meaningful, i.e. that they affect the variance; interventions on the mean, intuitively, are not informative as they shift the density graph by a constant, without affecting its shape (the gradient and the Hessian of the density, where information about the causal graph lies). ˜3 requires that the Jacobian of the inverse of the mixing function is informative about the causal structure at the mean of $\mathbf{S}$ (and it’s almost surely verified over $\mathbf{X}$ samples, under some generic regularity conditions on $\mathbf{f}$). The reason behind it is that we probe the identifiability of the Jacobian’s support at the mean. The only real simplifying constraint is ˜4 of the Gaussianity of the sources, which is, however, not new in the literature (see e.g.rolland2022score). Later, we discuss why this assumption is needed in the paper and potential ways to relax it.
@@ 4.1 Identifiability from second order derivatives of the log-likelihood
The assumption basically requires that there is sufficient variability between the different environments. Similar requirements of sufficient variability are ubiquitous in the nonlinear ICA literature (e.g. hyvarinen2016time; khemakhem2020icebeem; lachapelle22disentanglement). Intuitively speaking, ˜5 is satisfied when, for each of the two groups of environments ($1,...,e_{1}$ and $e_{1}+1,...,e_{1}+e_{2}$), each source $S_{j}$ is subject to rescaling. To see that, consider the LHS of the first equation: $\lambda_{j}^{i}=1$ for each $i=1,...,e_{1}$ corresponds to the case when the variable $S_{j}$ is never subject to rescaling in any of the environments, and indeed yields a violation of the assumption. Note that even if $S_{j}$ is subject to rescaling for some index $i$, the values of $(\lambda_{j}^{i})_{i\in[e_{1}]}$ can always be tuned such that the assumption is violated; however, this corresponds to pathological choices of the rescaling coefficients, which never occur in general (shown in Proposition˜3 in the appendix).
@@ 5 Empirical results
In this section, we report and analyse empirical results that validate our theory. Our experiments on synthetic data show that if the assumptions of Theorem˜1 hold, the causal direction can be recovered from the data. We focus on bivariate graphs, commonly adopted as the easiest yet non-trivial setting for testing identifiability (e.g., hoyer2008anm; zhang2009pnl; immer2022on).
@@ 5.1 Synthetic data generation
We generate synthetic data from bivariate causal models with independent noise terms, sampled from a normal distribution with unit mean and covariance entries uniformly drawn between $[1,1.5]$. Given the variables $x_{1},x_{2}$ and the graph $x_{1}\to x_{2}$ we consider the following causal mechanisms that comply with the assumptions of Theorem˜1: (i) $x_{2}:=s_{1}^{2}\operatorname{arctan}(s_{2})+s_{2}^{3}$ (ii) $x_{2}:=s_{1}^{2}s_{2}+\operatorname{arctan}(s_{2})$ (iii) $x_{2}:=s_{1}^{2}+\operatorname{arctan}(s_{1})s_{2}+s_{1}s_{2}^{3}$. Note that any of these models can not be reparametrized to a post nonlinear or location scale noise model, which are the most general SCMs identifiable from pure observations (zhang2009pnl; immer2022on). Additionally, we consider data from a linear Gaussian model, notably non-identifiable. We run experiments on datasets with $\{3,6,9\}$ environments. For each environment, we generate $2000$ observations. In Section˜C.4, we discuss experiments with non-Gaussian independent sources. Interestingly, these additional results seem to support our hypothesis that Theorem˜1 could be extended to other source distributions.
@@ 5.2 Analysis of the experimental results
The simplified pseudocode is found in Algorithm˜1 (a detailed version is presented in Section˜C.3). The steps in our procedure closely follow the proof of Theorem˜1: this approach to algorithmic design is not necessarily the best, which is why we highlight that our method is not within our main contributions. For a single inference, the input is the data tensor $\widehat{X}\in\mathbb{R}^{k\times n\times d}$: for each environment from $1$ to $k$ it consists of a dataset with $n$ observations of $d$ causal variables. Additionally, we are given the sets $I_{1},I_{2}\subset[k]$ of indices that split the auxiliary environments into two groups, as required by our theory. The first environment is taken as the base one. We have two steps where statistical estimation is involved: (i) For each environment, the gradient and the Hessian of the log-likelihood are approximated via the Stein gradient estimator, introduced in li2018gradient and popularized in causal discovery by rolland2022score; montagna23nogam; (ii) For each environment $i\in[k]$, we need to find the observation $j\in[n]$ such that $\mathbf{f}^{-1}(\widehat{X}[i,j])\approx\mu_{\mathbf{S}}$, that is, the data point generated mixing the source vector at the mean. Fortunately, this can be consistently estimated from the score $\nabla\log p_{\mathbf{x}}$, as we demonstrate in Proposition˜2 in the appendix. These two steps are achieved by Algorithm˜1 at the end of the first for loop. At this stage, all statistical quantities have been estimated: we note that, being the Stein estimator consistent, the algorithm is correct in the infinite sample limit. In the second for loop, we take the points at the estimated mean that we previously found, and compute the difference of the Hessians between the base and auxiliary environments: this exactly mirrors the first equality in Equations˜9 and 10 of Lemma˜1. Next, in the algorithm’s notation, we compute
@@
Then, we solve the linear system $\widehat{H}_{\textnormal{diffs}}[1]M=\widehat{H}_{\textnormal{diffs}}[2]$ to find $M$. In the infinite samples limit Equation˜13 is a precise equality, such that $M$ and $\Omega_{1}^{-1}\Omega_{2}$ are similar: diagonalizing $M$ we find $J_{\mathbf{f}^{-1}}$ up to a scaled permutation. The permutation indeterminacy is removed leveraging the assumption that the causal graph is acyclic via standard arguments (see shimizu2006alinear and reizinger2023jacobianbased). Finally, the algorithm returns the estimated support of the inverse Jacobian.
@@ A.4 Proof of Theorem 1
By Lemma 1 in reizinger2023jacobianbased, the permutation indeterminacy can be uniquely determined and thus removed. Given that by ˜3 the Jacobian of $J_{\mathbf{f}^{-1}}(\mathbf{x})$ is faithful to the causal graph, the claim is verified.
@@ Appendix B Independent component analysis
In this section, we present a primer on the problem of Independent Component Analysis (ICA), based on the content of Section 2 in buchholz2022function. ICA seeks to recover latent sources from their observed mixtures. We assume a hidden random vector $\mathbf{S}\in\mathbb{R}^{d}$ with independent coordinates and observations generated by
@@
In general (i.e., for $(\mathcal{F},\mathcal{P})$ arbitrarily large), the ICA problem is not identifiable for reasonable $\mathcal{S}$. Notable example comes from the Darmois construction or constructions based on measure-preserving transformations. Several results in the literature have studied which conditions on $(\mathcal{F},\mathcal{P})$ can help identifiability. Most notably, buchholz2022function shows that when $\mathcal{F}$ represents the class of conformal maps, identifiability is guaranteed up to trivial indeterminacies. If heterogeneous data are considered (e.g., in the multi-environment setting of this paper), identifiability was shown in the general case (hyvarinen2016time).